OpenStudy (anonymous):

If 2x + y = 5, what is the value of 4x?

OpenStudy (nurali):

2x + y = 5 2x=5-y x=5-y/2 4x 4(5-y/2) 20-4y/2 10-2y

OpenStudy (anonymous):

ah i see now, thank you!